# Zeta Zeros on the Critical Line

### v1

@@ 4. Montgomery’s Simple Zero Method if RH is not assumed
which proves, if $1\leq C<3/2$, that the proportion of simple zeros of $\zeta(s)$ on the critical line is $\geq 3-2C+o(1)$. ∎
@@ 5. Two Results on Critical Zeros
Assuming all of the zeros of $\zeta(s)$ with $T<\gamma\leq 2T$ lie in the region
@@
Then if $b=0.3185$, we have for the zeros of $\zeta(s)$ with $T<\gamma\leq 2T$ that at least 2/3 are simple, at least 2/3 are on the critical line, and at least 1/3 are both simple and on the critical line.
Can we get that asymptotically $100\%$ of the zeros are on the critical line by this method? Since the proportion of simple zeros is equal to that of critical zeros in Theorem 2, we need to use a conditional result on pair correlation that does not assume RH and implies that $100\%$ of the zeros are simple.
@@
$\sum_{\begin{subarray}{c}\rho,\rho^{\prime}\\ 0<\gamma,\gamma^{\prime}\leq T\\ 0<\gamma-\gamma^{\prime}\leq\frac{2\pi\lambda}{\log{T}}\end{subarray}}1\sim\left(\int_{0}^{\lambda}\left(1-\left(\frac{\sin{\pi u}}{\pi u}\right)^{2}\right)du\right)\frac{T}{2\pi}\log{T}.$
Recall that Montgomery’s Pair Correlation Conjecture (PCC) was the first suggestion that zeros of $L$-functions obey a random matrix eigenvalue model; for $\zeta(s)$, it is the Gaussian Unitary Ensemble (GUE) model.
Since PCC only concerns the non-zero vertical distance between zeros, it is unaffected by RH. On the other hand, nothing in PCC by itself implies directly that any zeros are on the critical line. Montgomery’s reasoning to support PCC used RH, but Gallagher and Mueller [GM78] proved that PCC implies asymptotically $100\%$ of the zeros are simple using a different method based on unconditional results of Selberg and Fujii. However, since RH was assumed and used in many parts of [GM78], we clarified this in [GLSS25a] to show the results on PCC hold without RH. As a consequence, by Theorem 2, we also obtain that asymptotically $100\%$ of the zeros are on the critical line.
@@
(5.2) $\sum_{\begin{subarray}{c}\rho,\rho^{\prime}\\ 0<\gamma,\gamma^{\prime}\leq T\\ |\gamma-\gamma^{\prime}|\leq\frac{2\pi\lambda}{\log{T}}\end{subarray}}1=\Big(1+o(1)\Big)\frac{T}{2\pi}\log T.$
Clearly (5.2) implies $C=1$, thus Theorem 2 follows from ES. This reduces the proof of Theorem 4 to proving without RH that PCC implies ES.
@@
The authors thank the American Institute of Mathematics for its hospitality and providing a pleasant environment where our work on this project began. The second author is also currently supported by the Inamori Research Grant 2024 and JSPS KAKENHI Grant Number 22K13895.

### v2

@@ 6. Simple and critical zeros using diagonal and symmetric diagonal terms
    the proportion of zeros of $\zeta(s)$ which are simple and on the critical line is $\geq 2-\mathbf{C}$,
  2. (ii)
    the average of the proportions of simple zeros and of critical zeros of $\zeta(s)$ is $\geq(3-\mathbf{C})/2$,
  3. (iii)
    the proportion of zeros of $\zeta(s)$ which are either simple or critical or both is $\geq(4-\mathbf{C})/3$.
Remark 1.
Notice that the proof of Theorem 2 actually only uses the assumption (6.1), and (i) of Theorem 3 already implies Theorem 2. The diagonal and symmetric diagonal terms are easy to analyze in terms of the properties of “simple” and “critical”, while we have dropped the non-symmetric horizontal terms which are more complicated. We return to this in the next section where we make use of “horizontal multiplicity”.
Remark 2.
Note that if $\mathbf{C}=4/3$ as in Theorem 1, then by (ii) of Theorem 3 the average of the proportions of simple zeros and critical zeros is asymptotically $\geq 5/6$. Thus at least one of these proportions is $\geq 5/6$, although the method doesn’t determine which. This agrees with the RH result in Theorem 1 where the proportion of simple zeros is $\geq 2/3$ and proportion of critical zeros is 1, thus giving an average proportion $\geq 5/6$. We obtain in (iii) of Theorem 3 that $\geq 8/9$ of the zeros are either simple or critical or both. Equivalently, the proportion of zeros that are multiple zeros off the critical line is $\leq 1/9$ .
Recall the multiset $\mathcal{Z(T)}$ from (3.1), and the set of distinct zeros $Z(T)=\{\rho:0<\gamma\leq T\}=\text{Supp}(\mathcal{Z}(T))$. We will consider $Z(T)$ the universe in what follows and take subsets $A\subset Z(T)$, and define the complement of $A^{\prime}$ in $Z(T)$ to be $A^{\prime}=\{\rho:\rho\notin A,\ 0<\gamma\leq T\}$. Similarly, we define the multiset $\mathcal{A}$ formed from $A$ by multiplicity and the complement $\mathcal{A^{\prime}}$ in the universe $\mathcal{Z(T)}$ by
(6.2) $\mathcal{A}=\{\rho\in A:0<\gamma\leq T,\ \rho\ \text{has}\ m_{\rho}\ \text{copies}\},\qquad\mathcal{A^{\prime}}=\{\rho\notin\mathcal{A}:0<\gamma\leq T\}.$
We will count the number of elements of $\mathcal{A}$ in two ways: $N(\mathcal{A})=|\mathcal{A}|$ where we count zeros with multiplicity, and $N^{*}(\mathcal{A})$ where we count zeros with multiplicity and also weighted by $m_{\rho}$. Thus
(6.3) $N(\mathcal{A}):=\sum_{\begin{subarray}{c}\rho\in\mathcal{A}\\ 0<\gamma\leq T\end{subarray}}1,\qquad N^{*}(\mathcal{A}):=\sum_{\begin{subarray}{c}\rho\in\mathcal{A}\\ 0<\gamma\leq T\end{subarray}}m_{\rho}.$
Note also that
$N(\mathcal{A^{\prime}})=N(\mathcal{Z})-N(\mathcal{A})=N(T)-N(\mathcal{A}).$
We will now work with the sets of simple zeros $S$ and critical zeros $C$ defined by
(6.4) $S:=\{\rho:\ 0<\gamma\leq T,\ \rho\ \text{simple}\ (m_{\rho}=1)\},\qquad C:=\{\rho:\ 0<\gamma\leq T,\ \rho\ \text{critical}\ (\beta=1/2)\}.$
We thus have the multisets $\mathcal{S}$ and $\mathcal{C}$ as defined in (6.2); note here $\mathcal{S}=S$. With this preparation, we are now ready to prove Theorem 3.
Proof of Theorem 3.
With the usual Venn diagram for the two sets $S$ and $C$, we partition the zeros in $Z(T)$ into the usual four disjoint sets according to being or not being simple/critical, form the corresponding multisets, and count with multiplicity to obtain
(6.5) $N(\mathcal{Z})=N(T)=N(\mathcal{S}\cap\mathcal{C})+N(\mathcal{S^{\prime}}\cap\mathcal{C})+N(\mathcal{S}\cap\mathcal{C^{\prime}})+N(\mathcal{S^{\prime}}\cap\mathcal{C^{\prime}}).$
The same decomposition in (6.1) gives on counting with multiplicity and weighting by multiplicity
(6.6) $N^{*}(\mathcal{S}\cap\mathcal{C})+N^{*}(\mathcal{S^{\prime}}\cap\mathcal{C})+2N^{*}(\mathcal{S}\cap\mathcal{C^{\prime}})+2N^{*}(\mathcal{S^{\prime}}\cap\mathcal{C^{\prime}})\leq(\mathbf{C}+o(1))N(T).$
Since $N^{*}(\mathcal{S}\cap\mathcal{C})=N(\mathcal{S}\cap\mathcal{C})$ and $N^{*}(\mathcal{S}\cap\mathcal{C^{\prime}})=N(\mathcal{S}\cap\mathcal{C^{\prime}})$, this equation becomes
$N(\mathcal{S}\cap\mathcal{C})+N^{*}(\mathcal{S^{\prime}}\cap\mathcal{C})+2N(\mathcal{S}\cap\mathcal{C^{\prime}})+2N^{*}(\mathcal{S^{\prime}}\cap\mathcal{C^{\prime}})\leq(\mathbf{C}+o(1))N(T).$
We eliminate the last two $N^{*}$’s in this equation using $N^{*}(\mathcal{S^{\prime}}\cap\mathcal{C})\geq 2N(\mathcal{S^{\prime}}\cap\mathcal{C})$ and $N^{*}(\mathcal{S^{\prime}}\cap\mathcal{C^{\prime}})\geq 2N(\mathcal{S^{\prime}}\cap\mathcal{C^{\prime}})$, where these are equalities if $\mathcal{S^{\prime}}$ only contains double zeros. Thus we obtain
(6.7) $N(\mathcal{S}\cap\mathcal{C})+2N(\mathcal{S^{\prime}}\cap\mathcal{C})+2N(\mathcal{S}\cap\mathcal{C^{\prime}})+4N(\mathcal{S^{\prime}}\cap\mathcal{C^{\prime}})\leq(\mathbf{C}+o(1))N(T).$
Now by (6.5) we see
$N(\mathcal{S^{\prime}}\cap\mathcal{C})+N(\mathcal{S}\cap\mathcal{C^{\prime}})+N(\mathcal{S^{\prime}}\cap\mathcal{C^{\prime}})=N(T)-N(\mathcal{S}\cap\mathcal{C}),$
and substituting this into (6.7), we obtain
$N(\mathcal{S}\cap\mathcal{C})+2(N(T)-N(\mathcal{S}\cap\mathcal{C}))+2N(\mathcal{S^{\prime}}\cap\mathcal{C^{\prime}})\leq(\mathbf{C}+o(1))N(T).$
Thus we obtain
(6.8) $N(\mathcal{S}\cap\mathcal{C})\geq(2-\mathbf{C}+o(1))N(T)+2N(\mathcal{S^{\prime}}\cap\mathcal{C^{\prime}}).$
From this we immediately have $N(\mathcal{S}\cap\mathcal{C})\geq(2-\mathbf{C}+o(1))N(T)$, which proves (i) of Theorem 3.
To prove (ii) of Theorem 3, we begin with the equalities
(6.9) $N(\mathcal{S}\cup\mathcal{C})=N(\mathcal{S})+N(\mathcal{C})-N(\mathcal{S}\cap\mathcal{C})\quad\text{and}\quad N(\mathcal{S}\cup\mathcal{C})=N(\mathcal{S}\cap\mathcal{C^{\prime}})+N(\mathcal{S}\cap\mathcal{C})+N(\mathcal{S^{\prime}}\cap\mathcal{C}),$
which on combining gives
$N(\mathcal{S})+N(\mathcal{C})=2N(\mathcal{S}\cap\mathcal{C})+N(\mathcal{S^{\prime}}\cap\mathcal{C})+N(\mathcal{S}\cap\mathcal{C^{\prime}}).$
On applying (6.5), this gives
$N(\mathcal{S})+N(\mathcal{C})=N(\mathcal{S}\cap\mathcal{C})+N(T)-N(\mathcal{S^{\prime}}\cap\mathcal{C^{\prime}}),$
and by (6.8) produces
$N(\mathcal{S})+N(\mathcal{C})\geq(3-\mathbf{C}+o(1))N(T)+N(\mathcal{S^{\prime}}\cap\mathcal{C^{\prime}})\geq(3-\mathbf{C}+o(1))N(T).$
To prove (iii) of Theorem 3, first note that by (6.7) and the second equation in (6.9) that
$2N(\mathcal{S}\cup\mathcal{C})-N(\mathcal{S}\cap\mathcal{C})+4N(\mathcal{S^{\prime}}\cap\mathcal{C^{\prime}})\leq(\mathbf{C}+o(1))N(T).$
Next, since $(\mathcal{S}\cup\mathcal{C})^{\prime}=\mathcal{S^{\prime}}\cap\mathcal{C^{\prime}}$, we have $N(\mathcal{S^{\prime}}\cap\mathcal{C^{\prime}})=N(T)-N(\mathcal{S}\cup\mathcal{C})$, and substituting this gives
$(4-\mathbf{C}+o(1))N(T)\leq 2N(\mathcal{S}\cup\mathcal{C})+N(\mathcal{S}\cap\mathcal{C})\leq 3N(\mathcal{S}\cup\mathcal{C}).$
7. Simple and critical zeros using horizontal multiplicity
Soundararajan pointed out to us an alternative approach for obtaining the results in the last two sections. He also showed us (iii) of Theorem 3.
In place of (5.2) and (5.4), we now use
(7.1) $\sum_{\begin{subarray}{c}\rho,\rho^{\prime}\in\mathcal{Z}(T)\\ \gamma=\gamma^{\prime}\end{subarray}}1=\sum_{\rho\in\mathcal{Z}(T)}H(\gamma)\leq\Big(\mathbf{C}+o(1)\Big)\frac{T}{2\pi}\log{T},\quad\text{where}\quad H(\gamma):=\sum_{\begin{subarray}{c}\rho^{\prime}\in\mathcal{Z}(T)\\ \gamma^{\prime}=\gamma\end{subarray}}1.$
If we assume RH, then $H(\gamma)=m_{\rho}=m_{\gamma}$ the usual multiplicity of the zero $\rho$. Without RH, then $H(\gamma)$ counts with multiplicity the number of zeros on the horizontal line $t=\gamma$. Thus $H(\gamma)$ is a “horizontal multiplicity” for the multiset of zeros on each horizontal line $t=\gamma$. Thus we see, for example:
  • If $H(\gamma)=1$, then on the line $t=\gamma$, we have one simple zero on the critical line.
  • If $H(\gamma)=2$, then on the line $t=\gamma$, we have either a double zero on the critical line or a pair of two symmetric simple zeros.
  • If $H(\gamma)=3$, then on the line $t=\gamma$, we have either a triple zero on the critical line, or a simple zero on the critical line and a pair of two symmetric simple zeros.
Note that to have at least a double zero off the critical line on the line $t=\gamma$, $H(\gamma)$ needs to be at least $4$.
Proof of Theorem 3.
Note $\mathcal{Z}(T)$ is partitioned into disjoint multisets of zeros by
$\mathcal{H}(k):=\{\rho\in\mathcal{Z}(T):H(\gamma)=k\},\qquad k\in\mathbb{N}.$
We also see from above that
$\mathcal{H}(1)\subset\mathcal{S}\cap\mathcal{C},\quad\mathcal{H}(2)\subset(\mathcal{S^{\prime}}\cap\mathcal{C})\cup(\mathcal{S}\cap\mathcal{C^{\prime}}),\quad\mathcal{H}(3)\subset(\mathcal{S}\cap\mathcal{C})\cup(\mathcal{S^{\prime}}\cap\mathcal{C})\cup(\mathcal{S}\cap\mathcal{C^{\prime}})=\mathcal{S}\cup\mathcal{C}.$
We now assume (7.1). If $H(\gamma)=1$, using an argument analogous to (4.5), we have
$N(\mathcal{S}\cap\mathcal{C})\geq\sum_{\begin{subarray}{c}\rho\in\mathcal{Z}(T)\\ H(\gamma)=1\end{subarray}}1\geq\sum_{\rho\in\mathcal{Z}(T)}(2-H(\gamma))\geq(2-\mathbf{C}+o(1))N(T),$
which proves (i) of Theorem 3. Next,
$\begin{split}(3-\mathbf{C}+o(1))N(T)&\leq\sum_{\rho\in\mathcal{Z}(T)}(3-H(\gamma))\leq 2\sum_{\begin{subarray}{c}\rho\in\mathcal{Z}(T)\\ H(\gamma)=1\end{subarray}}1+\sum_{\begin{subarray}{c}\rho\in\mathcal{Z}(T)\\ H(\gamma)=2\end{subarray}}1\\ &\leq 2N(\mathcal{S}\cap\mathcal{C})+N(\mathcal{S}^{\prime}\cap\mathcal{C})+N(\mathcal{S}\cap\mathcal{C}^{\prime})=N(\mathcal{S})+N(\mathcal{C}),\end{split}$
which proves (ii) of Theorem 3. For (iii) of Theorem 3, we have, since $\mathcal{H}(3)$ is disjoint from $\mathcal{H}(1)$ and $\mathcal{H}(2)$ and contained in $\mathcal{S}\cup\mathcal{C}$,
$\begin{split}(4-\mathbf{C}+o(1))N(T)&\leq\sum_{\rho\in\mathcal{Z}(T)}(4-H(\gamma))\leq 3\sum_{\begin{subarray}{c}\rho\in\mathcal{Z}(T)\\ H(\gamma)=1\end{subarray}}1+2\sum_{\begin{subarray}{c}\rho\in\mathcal{Z}(T)\\ H(\gamma)=2\end{subarray}}1+\sum_{\begin{subarray}{c}\rho\in\mathcal{Z}(T)\\ H(\gamma)=3\end{subarray}}1\\ &\leq 3N(\mathcal{S}\cap\mathcal{C})+2N(\mathcal{S}^{\prime}\cap\mathcal{C})+2N(\mathcal{S}\cap\mathcal{C}^{\prime})=2N(\mathcal{S}\cup\mathcal{C})+N(\mathcal{S}\cap\mathcal{C})\\ &\leq 3N(\mathcal{S}\cup\mathcal{C}).\end{split}$
@@ 8. Two Results on Critical Zeros
Assume that all of the zeros of $\zeta(s)$ with $T<\gamma\leq 2T$ lie in the region
@@
Then if $b=0.3185$, we have for the zeros of $\zeta(s)$ with $T<\gamma\leq 2T$ that at least 2/3 are simple and on the critical line.
We define the average spacing between consecutive zeros by
Thus in Theorem 4 we are assuming all the zeros are within a vertical distance of $b/(4\pi)$ times the average spacing between consecutive zeros at height $T$.
The following conjecture gives an asymptotic formula for the number of pairs of zeros counted with multiplicity within a distance $\lambda$ times the average spacing.
@@
$\sum_{\begin{subarray}{c}\rho,\rho^{\prime}\in\mathcal{Z}(T)\\ 0<\gamma-\gamma^{\prime}\leq\frac{2\pi\lambda}{\log{T}}\end{subarray}}1\sim\left(\int_{0}^{\lambda}\left(1-\left(\frac{\sin{\pi u}}{\pi u}\right)^{2}\right)du\right)\frac{T}{2\pi}\log{T}.$
Montgomery first made this conjecture in [Mon73]. From this conjecture sprung the discovery that the zeros of $L$-functions obey a random matrix eigenvalue model; for $\zeta(s)$ this is the Gaussian Unitary Ensemble (GUE) model.
Since PCC only concerns the non-zero vertical distance between zeros, it is unaffected by RH and nothing in PCC by itself implies directly that any zeros are on the critical line. Montgomery’s reasoning to support PCC used RH, but Gallagher and Mueller [GM78] proved that PCC implies asymptotically $100\%$ of the zeros are simple using a different method based on unconditional results of Selberg and Fujii. However, since RH was assumed and used in many parts of [GM78], we clarified this in [GLSS25a] to show the results on PCC hold without RH. As a consequence, by Theorem 2, we also obtain that asymptotically $100\%$ of the zeros are on the critical line.
@@
(8.3) $\sum_{\begin{subarray}{c}\rho,\rho^{\prime}\in\mathcal{Z}(T)\\ |\gamma-\gamma^{\prime}|\leq\frac{2\pi\lambda}{\log{T}}\end{subarray}}1=\Big(1+o(1)\Big)\frac{T}{2\pi}\log{T}.$
Clearly (8.3) implies $\mathbf{C}=1$, thus Theorems 2 and 3 follows from ES. This reduces the proof of Theorem 5 to proving without RH that PCC implies ES.
@@
We mention that Selberg [Sel92] and Bombieri and Hejhal [BH95] have both made use of ES in their work on linear combinations of $L$-functions. In particular, the latter paper made use of Montgomery’s Fejér kernel method used in Theorem 1 to obtain partial results without using ES.
@@
The authors thank the American Institute of Mathematics for its hospitality and providing a pleasant environment where our work on this project began. The second author is currently supported by the Inamori Research Grant 2024, JSPS KAKENHI Grant Number 22K13895, and Kyushu University International Research Leader Training Program (EBXU0101).